|NEWS up to 31 december 2000|
|16 Dec 2000||We have a problem with the radio link between the logger and the computers. Will be looked after next Monday...|
|11 Dec 2000||1. Logger restarted to begin with new shorter files. 7 days
graphs will be rebuilt during the next days.
2. We are in the test phase of our migration from macro-based file handling to a Linux solution based on perl scripts and gnuplot making png graphs This system is much more transparent than the baroque DOS solution running for the moment.
A test system is running and available under www.restena.lu/meteo_lcd/test/...
|29 Nov 2000||Lowest total ozone concentration of the year (220.4
DU)!; nothing to be afraid of: this is the season for
record low concentrations: see the KMI
There was a disruption in the daily graphs.. ok now (15:00 UTC)
|01 Nov 2000||Logger has to be restarted, lcd_uv.dat becoming too large.
This is the reason of faulty 7days graphs. The new graphs will recover
starting midnight. There was a power shortage yesterday (31Oct2000) from
09:30 to about 11:30.
The Sep00.dat and Oct00.dat files will be uploaded asap...
|18 Oct 2000||Transmission problem; some data missing on graphs, but present in lcd_uv data file (wait until tomorrow)|
|07 Oct 2000||Spurious transmission problems with one of the wireless modems... Not all data availbale on graphs for today.|
|01 Oct 2000||Small transmission problem today; will be corrected tomorrow...|
|26 Sep 2000||Restart of logger to correct 7 days graphs problem. Near-live graphs will be rebuilt during the next days|
|22 Sep 2000||1. Maintenance work on the ozone sensor and NOx sensors by Envitec
SA; recalibration of the latter.
2. There is a problem with the update of the 7 days graphs; the ASCII data are fine. Will look after asap...
|09 Sep 2000||Repair tried today at 15:30 to bring back the Peltier element to normal operation. For the moment, the element seems to be ok... but the problem could be an obscure recurrent one...|
|30 Aug 2000||The Peltier device which keeps the temperature of our main UVB sensor constant at approx. 25 degrees Celsius does not work anymore; the readings of the sensor are completely nonsens (check the positive UVB values at dark night!). This is a major malfunction, that can not be corrected in a short time. The plots of the UVB (mMED/h) irradiance and the UV index should be discarded.|
|19 Aug 2000||1. the directory with the .dat files was
unaccessible due to webserver change of our provider. Ok now
2. logger relaunched, near-live graphs will be rebuilt during the next days
|18 Aug 2000||I am back from a longer leave in the USA. The near-live graphs have stopped working after 05 Aug 2000, but the ASCII files contain the correct data. This problem will be solved asap...|
|13 Jul 2000||1. replacement pump mounted in CO2 sensor at 09:00
2. do not be afraid be extreme high NO and NO2 concentrations: an Diesel air compressor is working at 0.5m of our container, and the exhaust blows directly into the inlet of the NOx and O3 sensors. Watch the decrease of O3 when NO is high!
Construction works will continue for the next weeks; if [NO] is low, the air compres has been stopped (this happens every day at ca. 15:00)
|12 Jul 2000||The pump of the CO2 sensor behaves poorly and stops erratically from time to time. Tried to remedy this morning at 09:00. Will it help?|
|05 Jul 2000||1. added old overhauled digital anemometer to
ch.61; watch both wind-speed channels (ch3 and ch61) to look for a possible malfunction of
the analog sensor at ch3.
2. problem with bad graphs solved (the ESR sensor was not correctly configured the 3rd July)
3. upload of jun00.dat
Work on logger means some disruption in near online graphs; things will stabilize during the next hours and days.
|04 Jul 2000||Problem with the automatic graphs...Please disregard them and look at the ASCII data files for the moment. Remedial as soon as possible, probably tomorrow mporning...|
|03 Jul 2000||1. the barometer Air_Pres had a bad configuration
table: for the past 2 months, the readings are approx. 20 mbar to high. The table has been
corrected, and readings are now ok. This work means that the logger had to be
reinitialized, so all graphs will restart from fresh. Please disregard impossible high
spikes on graphs of today afternoon!
2. the dew-point signal (NTC_Dew) is much too high since a couple of weeks. There will be no fix to this problem in the near future
3. the anemometer Air_Velo readings seem good for small wind velocities, but too low for higher wind speeds. We will try to reinstall the old digital model at ch.61 during the next days, so check both channels..
|29 Jun 2000||Spooky house logo changed at home page: if everything is ok, you will see the friendly sun and green meadows logo, if there are problems, the spooky house will warn you!|
|27 Jun 2000||Mysterious intermittant breakdownof the CO2 sensor during the early morning hours; returns spontaneously to correct operation. Will be check if access to the container is free again (it is blocked for the moment by construction work going on...)|
|22 Jun 2000||Some configuration tests with the wireless modems during the after-noon; upload to the Internet was disabled for a couple of hours.|
|21 Jun 2000||..we are back on the air.
"Repaired" Proxim radio-modems have been returned to us...with a non working
configuration (how could it be different?). After much sweating in a 36°C container and
on a burning roof, we managed to build a working (slowly working to be precise)
configuration. All graphic files should be built up correctly during the next days.
Let me repeat how sorry the crew of meteo_lcd is on this radio-modem problem. We have only one source for this type of (very expensive) equipment, so we can not switch to something more reliable. In about 8 months, we will be able to connect the whole stuff directly to our LAN, so the radio-modem and all its related torture and suffering will then be history.... Again, thank you very much for your patience!
|15 Jun 2000||Upload of the May 2000 data. From 24th on, our new NO/NO2 sensor (AC31M from Environnement SA) is operational. There is some hope that new Nomadic wireless modems will be available next week.|
|19 May 2000||Upload to the Internet of the April 2000 data...
Many, many data are missing, due to disruption caused by our moving... do not rely on this
month to make you PHD in climatology!
A tentative radio modem repair was unsuccessful... these Nomadic modems (2.4 Ghz, DSSS type) from Proxim are not the jewel in the crown of the company. It is a pity that we do not have alternative communication means for the moment, and are forced to wait until we get a working couple of these devices!
|25 Apr 2000||Bad and good news:
The good ones: all sensors except NOx are again working in our new container. The old anemometer (ch.61, Air_Speed) has been replaced by a more precise analog model, whose data are now on ch.4 as "Air_Velo" (short for air velocitiy); ch.4 gave the soil temperature, which is not measured anymore (..well the soil always was a piece of concrete wall!)
The bad ones: one of the radio-modem broke down again... we have no replacement part on stock. Everything is done to hasten the repair!
|07 Apr 2000||The problem with the uncorrect graphs has been
solved...what works should now be graphed correctly; as always, resync will take some
The following sensors are not yet installed: CO2, Air_NTC, Dew_NTC, NO,NO2
These are not functioning correctly: Airspeed, Wind_Direction
The internal temperature (UVB_Tin) of the UVB_Biometer seems to be tto high (32 C instead of 25 C), but this could be a false alarm, as the sensors seems working correctly.
We put together a very tiny album with pictures on our moving out and in (photos are mostly under 30Kb size, and have not yet been beautified, due to lack of time). Please click here!
|06 Apr 2000||Solar and rain sensors now ok; I am unable to get the anemometer channel working...sorry, you will have to wait. The measured ozone data are correct (see the Ascii files), but the graphs are not (yet)|
|05 Apr 2000||Finished (well... almost!) moving into a new
container. We are exhausted, but are on the air again, albeit in an uncomplete manner:
here the sensors that do not yet work correctly or are not yet (re-)installed:
total solar radiation, wind speed, dew-point sensor, CO2, NO and NO2.
These are ok: air temp, air humidity (soil temp. will not be measured again, remember we are on a roof! The channel is reserved for future use), barometric pressure, O3, UVA,UVB, UVB_Bio, UVA_Bio. There will be some hickups before everything is in sync again.
I will be unavailable for about two weeks time. On my return, we will finish the job as fast as possible (we are waiting for a brand-new NOx sensor) and put some pictures on this site. Thanks for your patience....
|03 Apr 2000||We are moving to a new container; the whole equipment must be dismantled and rebuilt at the new facility. During this work (which may extend up to Thursday 6th April) there will be no live data!|
|29 Mar 2000||The radio modem has been replaced, and we are back on the air (still without O3 and NOx!). Many thanks to Mr. David Longree from 4i company for a quick help! As usual, synchronisation of 7 days graphs will take some time...|
|27 Mar 2000||A first try to repair the broken down wireless link failed; the modules are back to the supplier for repair. This will take between two and four weeks. Meanwhile meteo_lcd is unable to transmit near-live data to the internet.|
|22 Mar 2000||Full strike disaster: the radio-modem broke down yesterday night. As this device is the only communication link between the datalogger and the computers managing the Internet transfers, no new data will be available for the moment. We try to find an emergency replacement.. but it will take some time.|
|20 Mar 2000||The CO2 sensor behaves erratically... randomly shuts off ...|
|04 Mar 2000||1. Logger had problems with the 29th Feb 2000 (the last graphs were
somewhat prophetical!). This has now been corrected.
2. Some fool cut of the power supply cable for the loggers batteries, which as a consequence ran flat. No data lost for Feb 2000
3. February data uploaded to the Internet. Please remember that there no O3, NO & NO2 sensors operating for the moment.
4. Logger has been restarted at 08:00; the graphs will be correctly rebuilt during the next day(s).
|19 Feb 2000||Rough times for meteo_LCD: due to the somewhat
chaotic rebuilding schedule on our site, we are forced to remove the O3 and NOx sensors
for some time. Probably the whole facility will have to shut down for several weeks in the
near future. A temporary replacement solution is in the planning, but if the past has told
us a lesson, it is not to trust the future!
Please be patient and bear this misfortune with us. We will do our best to return to normal operation as soon as possible....
|06 Feb 2000||Again a problem with our main power supply, to to continous work on the building. We will try to draw a new power line as soon as possible....|
|29 Jan 2000||There is some disruption in service, due to a power outage on our main supply line. Logging has been restarted at 09:05, and graphs and files will be rebuilt during the next hours and days|
|26 Jan 2000||1999.zip file with all data from 1999 now available in subdirectory /data. All monthly files have been rechecked, missing data replaced by 000 to keep full file length.|
|25 Jan 2000||Back from the microbes! The Dec99 data have been transferred to the /data directory. The complete zip file with all 1999 data (1999.zip) will be available in the very near future, after some final checks.|
|16 Jan 2000||I'm laid down by a pneumonia for the next week... so the system will have to take care for itself. Dec 1999 data and full year 1999 zip file will be made when I am back...|
|11 Jan 2000||We had a small problem with Y2K, so graphs and data were not properly
sent do the Internet from 30Dec99 on. No data have been lost, as everything remained
stored in the logger, which does not store the year data and thus is immune to Y2K.
December 99 data will be transferred in a short time....
Logger is restarted today on 09:00 UTC
|20 Dec 1999||The two UV Biometers (recalibrated at Solar Light) are again on-line. The Delta_T UV sensors are on channels 7 and 8, as before. The output of the Delta_T UVB sensor is now given in mW/m2; please regard these readings with care, as the calibration factor is in a very basic alpha state!|
|10 Dec 1999||Logger halted and new configuration settings made for the Guardian II.
Logger restarted at 08:30.
Nov99 data uploaded to the web server. 7 days graphs will be rebuilt during the next days.
|09 Dec 1999||Gascard CO2 sensor replaced by recalibrated Guardian II (which is back
from the manufacturer). No CO2 data available up to 10 Dec 99 08:30.
Kroneis Dewpoint sensor removed for repair. Please disregard readings of channels 12 (KAI) and 13 (KDE) from now on!
|27 Nov 1999||To shorten the long global data file lcd_uv.dat, logger has been restarted (a new lcd_uv.dat start today at 11:00). All 7 days graphs will be rebuilt in the next days. The communication over a wireless Nomadic modem is not without problems: we try to get a smoother operation by avoiding the transfer of too large data files.|
|05 Nov 1999||Guardian II Co2 sensor definitively out of action (IR lamp blown). Will be replaced temporarly by Gascard sensor asap....A transceiver module in the optical line has been replaced, so we are back on the wire for the moment.|
|29 Oct 1999||We had a lot of problems with our CO2 sensor last week. Was replaced
temporarly by a Gascard sensor the 24th October; take CO2 readings of last 7 to 10 days
with caution.Today at 07:00 "old" Guardian was put again into action, after a
complete recalibration. There is no offset problem anymore.
We also changed our RS232 line form optic (there was an intermittent problem) to "old" copper, and today from copper to a wireless radio modem. Due to heavy construction work on our site we live in a state of permanent chaos and impeding disaster. To free the way for big lifting operations, we will remove the cables and lines and intend to stay wireless from today on...!
|13 Oct 1999||Oops... some wiring problem and a reconfiguration goof played havoc with the graphs and the UV data. Errors corrected this morning at 06:15. Good data will start at 06:30; good graphs will be rebuild automatically. Channels 7 and 8 have no sensors attached. Readings are fake data!|
|11 Oct 1999||Solar Light UV Biometers (channels 9 and 10) back to the manufacturer for
channels 9 & 10 now give the readings of the corresponding Delta_T sensors. Read here!
The 7 days graphs will rebuild during the next days.
|02 Sep 1999||Due to a nasty hardware failure, all data from the 9th Aug99 to 1st Sep99 are lost. Please share our sorrow....|
|Read a short report on the influence of the 11th Aug.99 solar eclipse on some meteorological data|
|Air humidity sensor bad starting 16:30. Dew-point signals (NTC_Air and NTC_Dew)also bad since some time. Remedial action will be taken during September 99. Sorry...|
|All data cleared from logger to enable a fresh restart. No data are lost! Seven days graphs will be rebuilt during the next days...|
|The Microtops II sensor is back from Solar Light (no compliments for the speedy work!) We just wait for a sunny day to resume measurements.|
|30 May 1999||The high spikes in the NO signal are NOT due to a sensor malfunction: heavy construction work is going on at the site of our station, involving useage of a lot of diesel machinery. This work happens from Monday to Saturday: so watch the calm down to "normal" values during the Sundays! Click here for a proof!|
|25 May 1999||1. Definitive recalibration of the NO/NO2 sensor
2. The filter of the rain sensor was choked up...replaced by a kludge. Be careful in interpreting rainfall readings of last weeks!
|12 May 1999||Maintenance work by Envitec SA on the NO/NO2 and O3 sensor. Recalibration of the NO/NO2|
|03 Apr 1999||..well, as for an offset of the CO2 sensor: this really was a BIG one. Actually the very low (-1600ppm) readings were due to a wiring error! Sensor works now fine again, but please DO NOT FORGET TO ADD +26PPM to the readings!|
|Original Guardian II CO2 sensor, refitted with a new pump and recalibrated is back into action at 16:00. Beware: the instrument has an uncorrectable zero offset of -26ppm. Please add +26ppm to all raw data of the ASCII files. This warning will be repeated in every header of the upcoming monthly data files. The automatically made graphs will be adjusted asap....When done, this will be reported here.|
|The Microtops II sunphotometer used for measuring total ozone is back to Solar Light Co. for recalibration. Sorry, no Dobson data available for the moment....|
|A new version of the report comparing the total ozone mesurements of the Microtops II with those done by a Brewer spectrometer at the RMI in Uccle; this time all available data from 1997 and 1998 are used.|
|Feb 1999 04||A new paper by Francis Massen et al. on "Cave volume computed on the behaviour of a blowing well..." in TERRA NOVA, vol.10, 131-135, 1998. Click here for abstract!|
|Jan 1999 31||Microtops II Dobson data for 1998 now available in data directory|
|Jan 1999 20||Upload December 1998 data and zipped file 1998.zip with all data from 1998|
|Dec 1998 14||Maintenance work on O3 and NO/NO2 sensors. Recalibration of the NO/NO2 sensor|
|Oct 1998 31||Maintenance work at the NO2/NO and O3 sensors. CO2
Guardian II sensor readings are dubious (my be too high). Replaced temporarly by a Gascard
NDIR sensor: be aware: this replacement sensor has not been recalibrated, so take its
readings with a pinch of salt...
Download area: "huuuge lcd_uv.dat" file stops for unknown reason several days too early. 7days.dat file is ok. This will be mended asap...
|Oct 1998 27||Recalibration of NO2/NO sensor (span adjusted with gas of 412 ppb NO concentration: this is the reason for the peak!)|
|Oct 1998 21||Major overhaul of NO2/NO sensor from 14:00 to 17:00. Please do not expect reliable data before some 3 to 4 days (time needed to stabilize instrument after changing the molybdene reaction chamber)|
|Oct 1998 19||Maintenance work on NO2/NO sensor at 09:30 (spike!)|
|Oct 1998 15||A report on radon measurements done in an underground gallery situated under the main reservoir of the pumping storage facility at Vianden, Luxembourg (authors: Francis Massen, Antoine Kies).. The report is written in German!|
|Oct 1998 02||NO2/NO sensor repaired. A Z80 Pio lost one of its legs! Sensor again functional from 14:00 on. Thanks to Envitec NV from Gent for quick assistance!|
|Sep 1998 20||Breakdown of NO2/NO sensor (processor malfunction) at 14:00|
|Sep 1998 10||There were several malfunctions during the last days:
1. Air Humidity sensor (#3) went bad from 01Sep98 15:00 on. Has been replaced 10Sep98 10:30
2. NO2/NO sensor malfunction due to intermittant bad conatct. Repaired 10Sep98 10:30
3. Seven days graphs stuck on 30 Aug98; the cause was an exceeding long data file. Has been corrected 10Sep98 10:30
4. Logger has been relaunched 10Sep98 07:30, which is now the start point of the new lcd_uv.dat file. Seven days graph will be built from this date on.
|Aug 1998 14||Operator of station back for general check: everything is ok... Download jul98.dat file. See partial data file aug98_1.dat for high ozone peaks around 10 Aug98 . These temporary peaks do not give me sleepless nights.|
|July 1998 11||The operator of this station will be unavailable for 4 weeks: that means that we all must have full confiance in the automatically running of the system! Minor maintenance, as filter replacement etc.. will be made as usual every 14 days. So dear friends, cross your fingers, hold your breath.. and see us again in a month!|
|July 1998 10||Maintenance work on NO2/NO sensor at 08:00. This explains peak
in NO2/NO graphs.
Heavy construction work with a lot of diesel fueled machinery has started since approx. a week. The exhaust fumes of the machines running 25m below the station seem to cause the NO peaks at hours different from the usual morning traffic peak at 06:00
|June 1998 12||Maintenance work on NO2/NO sensor at 08:00. This explains peak in NO2/NO graphs.|
|May 1998 29||Maintenance work on NO2/NO sensor at 09:00. This explains down-peak in graphs.|
|May 1998 22||New NO2/NO
sensor starts working at 16:00. Graphical data will be available asap. This
installation and maintenance work done on the ozone sensor means the 7 days graphs will
restart the 23 May. Down peaks are due to instruments being powered down.
NO2 and NO are on channels 16 and 17; the UVB and UVA Biometers internal Peltier temperatures are moved to channels 18 and 19. These two signals are again available starting 16:00
|May 1998 18||Installation of a new NO2/NO sensor in progress. This explains the down-peak in ozone of today. When sensor works, this will be announced here. Meanwhile the data of the internal UV_Biometers temperature sondes are removed until new sensor is working. The NO2 and NO data will be on channel 16 and 17, the UVB_Biometer internal temperatures are being moved to channels 18 and 19.|
|May 1998 09||Logger system hangup at 18:00. No data available from 09May 18:30 to 10May 12:00. Probably some temporary problems before graphs are again correctly synchronized.|
|May 1998 06||CO2 sensor changed against same model (Gasgard II)|
|April 1998 10||Logger system hangup. Not many data are lost (a few hours). System restarted at 16:30. All graphs will restart at that point...|
|March 1998 24||We had a severe malfunction on our main Internet computer. No near-live data were available for 3 days. No data are lost, as everything continued to be stored in the logger. We are on the air again... 16:40 UTC|
|March 1998 05||All detailed data from the Mircotops II total ozone measurements can now be found in the data area under the name of mtop0398.dat|
|February 1998, 16||Logger crash between 10:00 and 11:00 approx. Restarted Feb. 17th at 08:30. The 7 days graphs are rebuilt starting 16 Feb 98. The cause of these two crashes might be a strong elm pulse of unknown origin.|
|February 1998, 11||Logger crash at 10:00 approx. Restarted at 13:30. The 7 days graphs are rebuilt starting 11 Feb 98.|
|February 1998, 09||Air temperature and air humidity sensors replaced by new ones.
Attention: Dewpoint sensor values seem much to high! Do not rely on them for the moment...
|January 1998, 23||Now graphs of 7 last days, including today! Also added separate graph of the UVI. Major reworking still going on... please be patient and do not count the remaining bugs!|
|January 1998, 11||A new report and comparison of personal UVB and UVI sensors!|
|January 1998, 05||Logger malfunction repaired; restarted at 10:00 UTC|
|January 1998, 05||All 1997 data available as 1997.zip file|
|October 1997, 15||New book by Francis Massen on the Climate of Moestroff Cave . This is definitively underground work!|
|October 1997, 01||New paper comparing the total ozone mesurements of the Microtops II with those done by a Brewer spectrometer at the RMI in Uccle|
|July 1997, 19||No more total ozone measurements at the Diekirch Site until September 2th
1997! Ozone checking will continue at another location (in the Provence, France), but
these measurements will not be published.
Comparing the total O3 measurements done until now at Diekirch with those of the Brewer instrument at Uccle (Belgium) shows extremely good concordance!
|July 1997, 17||Air Humidity sensor definitively out of order. Will be replaced only at the end of Sept.97. Please use dew point sensor readings!|
|July 1997, 09||Overhaul and recalibration of the ozone sensor by Envitec SA (B-Gent)|
|June 1997, 08||New Paper on UV
related health aspects for the general public; by Nico Harpes. In german!
|May 1997, 24||Foire Internationale de Luxembourg 24 May to 1 June 97:
Meteo_LCD is in Hall 3D, boot 22 (DEFI)
|May 1997, 16||A UVI-Poster explaining (confusing?) units and definitions|
|April 1997, 14||Explications on the UVI bulletins|
|March 1997, 18||Start of ISDN data transmission to Internet provider Restena|
|March 1997, 13||New continuous working CO2 sensor (Guardian II) on channel
Internal temperature of thermally stabilized UVB_Bio and UVA_Bio sensors at channels 16 and 17. Should always be close to 25 Celsius.
Sorry for any disruption caused today by upgrade-work!
|March 1997, 09||Total Ozone spot measurements available now!|
|March 1997||UVB Biometer back from calibration; restart working 06 March 97 08:30|
|February 1997||Meteo_LCD participates in German UV-dosimetry|
|Back to Home-Page of Meteo_LCD|